### v1

# Robust Distributed Gradient Descent to Corruption over Noisy Channels

Abstract
Distributed gradient descent algorithms have come to the fore in modern machine learning, especially in parallelizing the handling of large datasets that are distributed across several workers. However, scant attention has been paid to analyzing the behavior of distributed gradient descent algorithms in the presence of adversarial corruptions instead of random noise. In this paper, we formulate a novel problem in which adversarial corruptions are present in a distributed learning system. We show how to use ideas from (lazy) mirror descent to design a corruption-tolerant distributed optimization algorithm. Extensive convergence analysis for (strongly) convex loss functions is provided for different choices of the stepsize. We carefully optimize the stepsize schedule to accelerate the convergence of the algorithm, while at the same time amortizing the effect of the corruption over time. Experiments based on linear regression, support vector classification, and softmax classification on the MNIST dataset corroborate our theoretical findings.
Index Terms:
Distributed gradient descent, Mirror descent, Adversarial corruptions, Noisy channels, Convergence rates
I Introduction
Driven by the growing scale of datasets in various learning tasks, distributed gradient descent has become increasingly important in modern machine learning  [2, 3, 4]. In this paradigm, there is a central parameter server coordinating information among all workers and each worker in the system processes partial gradients using its subset of data [5, 6]. These partial gradients are then collectively used to update the global model parameters. Such a distributed framework not only expedites the learning process but also manages datasets that are too large to be handled in the memory of a single worker or server.
However, there are significant challenges in the use of distributed gradient descent such as the communication overhead [7], the necessity for resilience against system failures [8, 9], and the presence of adversarial attacks [10, 11, 12, 13]. In a distributed framework, workers may become unreliable (called Byzantine workers in the sequel), which can disrupt the learning process. For instance, as demonstrated in [14], the presence of even a single Byzantine worker can undermine the convergence of a distributed gradient descent algorithm.
In this paper, we consider a new problem concerning the distributed gradient descent algorithm but we take into account a novel and critical aspect that has, to the best of our knowledge, not been considered before. We consider the presence of adversarial corruptions at each worker in addition to the usual random (Gaussian) noises in the uplink and downlink channels. The corruptions and noises affect the accuracy of partial gradients and parameters. Each worker experiences adversarial corruption, with the total corruption budget being bounded over the entire time horizon. Our convergence analysis of a modified version of a (lazy) mirror descent-type algorithm addresses the worst-case scenario of these corruptions subject to a long-term constraint on the total corruption budget. We show both theoretically and empirically that our algorithm RDGD and its variants such as RDGD-Restart  can be suitably tuned to be corruption tolerant.
I-A Main Contributions
Our main contributions are as follows:
  1. 1.
    We formulate the problem of distributed gradient descent with adversarial corruptions on the partial gradient vectors held by each of the $m$ workers. We impose a bound $C(T)$ on the total amount of corruptions over the $m$ workers and over the time horizon $T$. The uplink and downlink channels are also subject to the usual Gaussian noises.
  2. 2.
    We design a mirror descent-inspired distributed gradient descent algorithm RDGD that is robust to corruptions. For smooth and convex functions, we show that the expected suboptimality gap of RDGD  scales as $O\big{(}\frac{C(T)}{m\sqrt{T}}\big{)}$.
  3. 3.
    We extend RDGD to the case of loss functions that are strongly convex in addition to being smooth. We bound the suboptimality gap of the enhanced algorithm RDGD-SC. The convergence behavior of RDGD-SC depends on different choices of the stepsize. For a certain choice of stepsize, we observe that the expected suboptimality gap decays as $O\big{(}\rho^{t}+\frac{C(t)}{m}\big{)}$ where $\rho\in(0,1)$ is a contraction factor and $t$ is number of iterations, while for another choice, the gap decays as $O\big{(}\frac{1}{t^{2}}+\frac{C(t)}{m\sqrt{t}}\big{)}$. Thus for the former, there is an exponentially decaying term and a term that does not cause the corruption to vanish; for the latter, there is a (slower) polynomially decaying term and a term that amortizes the corruption.
  4. 4.
    The above observation motivates us to design a hybrid algorithm RDGD-Restart  that exploits the benefits of each choice of the stepsize. RDGD-Restart switches the stepsize schedule at an analytically determined transition time. In RDGD-Restart, we first exploit the exponential decrease of the suboptimality gap, and eventually amortize the corruption by changing the stepsize.
  5. 5.
    The proposed algorithms are assessed numerically through several canonical machine learning models such as linear regression, support vector classification, and softmax classification (on the MNIST dataset). All experimental results corroborate our theoretical findings.
I-B Related Works
There are many studies that analyze the effect of random corruptions in (stochastic) gradient descent [15, 16, 17] and its accelerated versions [15, 18, 19]. Our study adopts a distinctly different approach in which we focus on the effect of malicious adversarial corruptions, instead of random ones, on distributed gradient descent-type algorithms. In this scenario, some workers send maliciously corrupted gradient information to adversely affect the system’s performance. In [20], the authors investigate the robustness of gradient descent in the face of the worst case of corruption, where adversaries can manipulate the corruption up to a certain bounded extent throughout the gradient descent process. In contrast to our distributed gradient descent formulation, [20] considers a centralized gradient descent scheme. Furthermore, the approach of [20] does not modify the traditional gradient descent algorithm to further mitigate the adversarial corruption’s effect.
Another line of research that relevant to our work pertains to Byzantine distributed gradient descent. Byzantine faults refer to the scenario where some nodes in the distributed network act arbitrarily or maliciously, potentially leading to incorrect or suboptimal outcomes. One line of solutions is based on the robust aggregation rules, including trimmed mean [12], coordinate-wise median [12], geometric median [11], and Krum [14]. However, these robust aggregation rules work under the assumption that the number of Byzantine workers is less than half the total number of workers. Our research approach bifurcates from this line of research since we allow the adversary to arbitrarily allocate a bounded amount of corruptions among all workers.
In addition, these robust aggregation methods can, in general, only recover the parameters obtained from vanilla gradient descent with high probability but not with probability one. To address this challenge, an algorithm called DRACO is presented in [10]. DRACO is a framework that applies repetition codes to the computational results of the workers. This requires roughly doubling the replication of each gradient computation, i.e., $12s+$ 1 replications of each partial gradient for $s$ malicious workers. Recently, [21] reduced the number of replications from $12s+$ 1 to $s+1$ by introducing an additional step to first detect the Byzantine workers. However, if the channels used for transmitting the parameters are noisy, this idea of replicating each partial gradient does not work perfectly well since the parameter server may fail to reliably select the majority among all the corrupted partial gradients.
The works on adversarial optimization are also closely related to some lines of work in the multi-armed bandit literature, particularly in the adversarial (and not stochastic) setting [22, 23, 24, 25]. In these settings, adversarial corruption is added to the random rewards, resulting in the agent receiving corrupted versions of the rewards. These approaches typically assume the adversary is capable of arbitrarily allocating corruption in a malicious manner within a total corruption budget. They then use various randomized policies to mitigate the effect of the corruption. However, our study focuses on distributed gradient descent, which is a first-order optimization method that is rather different scenarios with bandit feedback.
I-C Outline
The rest of this paper is organized as follows. The framework of distributed gradient descent in the presence of adversarial corruptions over noisy channels is presented precisely in Section II. The proposed RDGD algorithm and the analysis of its expected suboptimality gap for smooth loss functions are developed in Section III. In Section IV, we additionally assume the loss function is strongly convex, design an extended algorithm RDGD-SC, and prove a bound on its expected suboptimality gap. The algorithm RDGD-Restart to mitigate corruption accumulation through a judicious choice of stepsizes is presented in Section V. Numerical experiments are presented in Section VI. Finally, Section VII concludes the paper by suggesting several avenues for further research.
II Problem Setting
In this section, we present the framework of distributed gradient descent in the presence of corruptions and noise in channels in both the downlink and the uplink. In addition, we present some basic notations and preliminaries that will facilitate the subsequent analysis.
II-A Distributed Gradient Descent
We assume that the samples in a dataset $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1}^{N}$ are drawn from an unknown distribution $\mathcal{D}$. Here $x_{i}\in\mathbb{R}^{p}$ is the $i$-th feature vector and $y_{i}\in\mathbb{R}$ is the corresponding scalar response (which could be discrete or continuous). Let $\Theta\subset\mathbb{R}^{p}$ be the set of model parameters. We seek to minimize the average of loss functions $L:\mathbb{R}^{p}\times\mathbb{R}\times\Theta\rightarrow\mathbb{R}$ each evaluated over a sample in $\mathcal{Z}$, i.e., we consider the problem
$\min_{\theta\in\Theta}\Big{\{}L(\theta)\coloneq\frac{1}{N}\sum_{i=1}^{N}L(x_{i},y_{i};\theta)\Big{\}}.$
In a centralized gradient descent scheme, the model parameter $\theta_{t}\in\Theta$ is updated at iteration $t$ as follows,
$\theta_{t+1}=\theta_{t}-\eta_{t}g_{t},$
where $g_{t}:=\frac{1}{N}\sum_{i=1}^{N}\nabla L(x_{i},y_{i};\theta_{t})$ is the full gradient ($\nabla L(x_{i},y_{i};\theta)$ is the gradient vector of $L(x_{i},y_{i};\theta)$ w.r.t. $\theta$) and $\eta_{t}$ is an appropriately chosen stepsize at iteration $t$.
When $N$ is large, it is inefficient to compute the full gradient $g_{t}$. Rather, one typically parallelizes the computation of $g_{t}$ by distributing the data points among $m\ll N$ workers. We assume the $N$ data samples in $\mathcal{Z}$ are uniformly partitioned among the $m$ workers. [^1] Without loss of generality, we assume $N$ is a multiple of $m$. In particular, the $i$-th worker maintains a subset $\mathcal{Z}_{i}$ of the dataset $\mathcal{Z}$, where $\mathcal{Z}_{i}\cap\mathcal{Z}_{j}=\emptyset$, $|\mathcal{Z}_{i}|=N/m$. Then, worker $i$ computes its partial gradient vector
$g_{i,t}:=\frac{1}{|\mathcal{Z}_{i}|}\sum_{(x,y)\in\mathcal{Z}_{i}}\nabla L(x,y;\theta_{t}).$
This partial gradient vector is then sent to the server.
II-B Noisy channel transmission
At the $t$-th iteration of distributed gradient descent, the parameter server broadcasts the current model parameter $\theta_{t}\in\Theta$ to all $m$ workers through noisy downlink channels. Worker $i$ receives a noisy version of $\theta_{t}$ as:
$\theta_{i,t}=\theta_{t}+v_{t}^{(i)},$ (1)
where $v_{t}^{(i)}=[v_{t,1}^{(i)},\dotsc,v_{t,p}^{(i)}]^{\top}\in\mathbb{R}^{p}$ is the $p$-dimensional noise vector at worker $i$ and time $t$. Furthermore, $v_{t}^{(i)}$ is a zero-mean random vector consisting of independent and identically distributed (i.i.d.) elements with variance $\mathbb{E}[(v_{t,k}^{(i)})^{2}]=\sigma_{t}^{2}$. Thus,
$\mathbb{E}[\|v_{t}^{(i)}\|_{2}^{2}]=p\sigma_{t}^{2},\quad\forall i\in[m].$
Now, the noisy partial gradient vector at worker $i$ for based on its local dataset $\mathcal{Z}_{i}$ can be written as
$g^{\prime}_{i,t}:=\frac{1}{|\mathcal{Z}_{i}|}\sum_{(x,y)\in\mathcal{Z}_{i}}\nabla L(x,y;\theta_{i,t}),$ (2)
where $\theta_{i,t}$ is the noisy parameter vector in (1).
II-C Corrupted gradients
At iteration $t$, instead of transmitting the true local gradients, due to Byzantine faults, each worker also suffers an arbitrary corruption to its partial gradient vector as follows:
$\bar{g}_{i,t}=g^{\prime}_{i,t}+\varepsilon_{i,t}.$ (3)
Here the corruption $\varepsilon_{i,t}$ is a vector in $\mathbb{R}^{p}$. Define the $\ell_{2}$ norm of the total corruption at iteration $t$ as:
$c_{t}=\bigg{\|}\sum_{i=1}^{m}\varepsilon_{i,t}\bigg{\|}_{2}.$ (4)
This quantifies the aggregate corruption in gradients, introduced by the adversary at iteration $t$. We denote the corruption vector as $c=[c_{1},c_{2},\dotsc,c_{T}]^{\top}\in\mathbb{R}^{T}$ and assume that there is an overall $\ell_{2}$ bound on it across the entire time horizon. The form of the bound we employ is
$\|c\|_{2}=\Big{(}\sum_{t=1}^{T}c_{t}^{2}\Big{)}^{1/2}\leqslant C(T).$ (5)
We emphasize that terms $\{\varepsilon_{i,t}\}$ and hence, $\{c_{t}\}$, and $c$ are deterministic quantities induced by Byzantine faults. These are in contrast to the stochastic noises in the (downlink and uplink) channels. Our subsequent analysis is thus applicable to the worst-case instantiation of these corruptions subject to the constraint in (5).
The $i$-th worker sends $\bar{g}_{i,t}$ defined in (3) to the parameter server. However, due to the presence of noise in the uplink communication channels, the parameter server receives the following perturbed version of $\bar{g}_{i,t}$ from worker $i$:
$\tilde{g}_{i,t}=\bar{g}_{i,t}+w_{t}^{(i)},$ (6)
where $w_{i,t}^{(i)}=[w_{t,1}^{(i)},\dotsc,w_{t,p}^{(i)}]^{\top}\in\mathbb{R}^{p}$ is a $p$-dimensional noise vector at worker $i$ and time $t$. Furthermore, $w_{t}^{(i)}$ is a zero-mean random vector consisting of i.i.d. elements with variance $\mathbb{E}[(w_{t,k}^{(i)})^{2}]=\sigma_{t}^{2}$. As a result,
$\mathbb{E}[\|w_{t}^{(i)}\|_{2}^{2}]=p\sigma_{t}^{2},\quad\forall\,i\in[m].$
An illustration of our distributed gradient descent setting with corruptions over noisy channels is presented in Fig. 1.
[…]
Figure 1: Illustration of the distributed gradient descent setting with corruption over noisy channels. The variables indicated in blue are random Gaussian noises, defined in (1) and (6). The variables indicated in red are deterministic adversarial corruptions specified in (3) and subject to the bound in (5).
II-D Preliminaries
Consider a continuously differentiable function $f:\Theta\to\mathbb{R}$, where $\Theta\subseteq\mathbb{R}^{p}$ is a closed and convex set. The convexity of $f$ is characterized by the following inequality:
$f(\theta^{\prime})\geqslant f(\theta)+\langle\nabla f(\theta),\theta^{\prime}-\theta\rangle,\quad\forall\,\theta^{\prime},\theta\in\Theta.$
Definition 1 (Smoothness)
A differentiable function $f:\Theta\to\mathbb{R}$ is $M$-smooth on $\Theta$ if for all $\theta,\hat{\theta}\in\Theta$,
$f(\hat{\theta})\leqslant f(\theta)+\langle\nabla f(\theta),\hat{\theta}-\theta\rangle+\frac{M}{2}\|\hat{\theta}-\theta\|_{2}^{2}.$
Definition 2 (Strong Convexity)
A differentiable function $f:\Theta\to\mathbb{R}$ is $\mu$-strongly convex on $\Theta$ if for all $\theta,\hat{\theta}\in\Theta$,
$f(\hat{\theta})\geqslant f(\theta)+\langle\nabla f(\theta),\hat{\theta}-\theta\rangle+\frac{\mu}{2}\|\hat{\theta}-\theta\|_{2}^{2}.$
Definition 3 (Convex Conjugate)
The convex conjugate of a function $\psi:\Theta\subset\mathbb{R}^{p}\to\mathbb{R}$, denoted as $\psi^{*}:\mathbb{R}^{p}\rightarrow\mathbb{R}\cup\{\pm\infty\}$
$\psi^{*}(z)=\max_{\theta\in\Theta}\{\langle z,\theta\rangle-\psi(\theta)\}.$
In the following, we assume that whenever we have to compute the convex conjugate $\psi^{*}$ of a function $\psi:\Theta\to\mathbb{R}$, this can be done efficiently and potentially in closed form.
The following is a corollary of Danskin’s Theorem [26].
Proposition 1
Let $\psi:\Theta\to\mathbb{R}$ be a differentiable and strongly convex function. Then, the gradient of its convex conjugate $\psi^{*}$ at any point $z$ is given by:
$\nabla\psi^{*}(z)=\operatorname*{arg\,max}_{\theta\in\Theta}\{\langle z,\theta\rangle-\psi(\theta)\}.$
Definition 4 (Bregman Divergence)
Let $\psi:\Theta\to\mathbb{R}$ be a continuously differentiable and strictly convex function. The Bregman divergence $D_{\psi}:\Theta^{2}\rightarrow\mathbb{R}$ is defined as:
$D_{\psi}(\theta,\theta^{\prime})\coloneq\psi(\theta)-\psi(\theta^{\prime})-\langle\nabla\psi(\theta^{\prime}),\theta-\theta^{\prime}\rangle.$
The Bregman divergence $D_{\psi}(\theta,\theta^{\prime})$ quantities the gap between $\psi(\theta)$ and its first-order approximation at $\theta^{\prime}$.
III Robust distributed gradient descent to corruption over noisy channel
In this section, we present Robust Distributed Gradient Descent (RDGD), an algorithm that is robust to corruptions over noisy channels. For the minimization of smooth loss functions, the algorithm is described in Algorithm 1. We also provide a bound on the expected suboptimality gap of RDGD.
III-A The RDGD  Algorithm
Algorithm 1 Distributed Gradient Descent Algorithm Robust to Corruptions over Noisy Channels (RDGD)
1:Initialization: Parameter vector $\theta_{0}\in\Theta$, dual vector $z_{0}=\nabla\psi(\theta_{0})$ for some $\mu$-strongly convex function $\psi$, algorithm parameters $\eta_{t}$, and $T$. $\theta_{1}=\theta_{0}$ and $\hat{\theta}_{0}=\theta_{0}$.
2:for $t=1,2,\ldots,T$ do
3:   Parameter server: Send $\theta_{t}$ to all $m$ workers over the downlink noisy channel.
4:   for all $i\in[m]$ do in parallel
5:      Worker $i$:
6:      Receive $\theta_{i,t}=\theta_{t}+v_{t}^{(i)}$ and calculate the corrupted local gradient $g^{\prime}_{i,t}$ in (2).
7:      Send $\bar{g}_{i,t}=g^{\prime}_{i,t}+\varepsilon_{i,t}$ to the parameter server via the noisy uplink channel.
8:   end for
9:   Parameter server:
10:   Receive $\tilde{g}_{i,t}=\bar{g}_{i,t}+w_{t}^{(i)}$ for all $i\in[m]$.
11:   Compute mean gradient $\tilde{g}_{t}\leftarrow\frac{1}{m}\sum_{i=1}^{m}\tilde{g}_{i,t}$
12:   Update model parameters:
13:   $z_{t}=z_{t-1}-\eta_{t}\tilde{g}_{t}$, i.e., $z_{t}=-\sum_{k=1}^{t}\eta_{k}\tilde{g}_{k}+z_{0}$.
14:   $\hat{\theta}_{t}=\frac{H_{t-1}}{H_{t}}\hat{\theta}_{t-1}+\frac{\eta_{t}}{H_{t}}\theta_{t}$, where $H_{t}=\sum_{k=1}^{t}\eta_{k}$.
15:   $\theta_{t+1}=\operatorname*{arg\,min}_{u\in\Theta}\{\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k},u-\theta_{k}\rangle+D_{\psi}(u,\theta_{0})\}$.
16:end for
17:Output: The parameter vector $\hat{\theta}_{T}$.
[…]
Figure 2: Illustration of primal and dual updates in RDGD.
At a high level, we utilize a modified lazy mirror descent approach with primal and dual updates. To initialize, we map the initial point $\theta_{0}$ to $z_{0}=\nabla\psi(\theta_{0})$ in the dual space via the mirror map $\nabla\psi$ for some suitably chosen $\mu$-strongly convex function $\psi$. At the $t$-th iteration of RDGD, the parameter server receives the corrupted partial gradients $\tilde{g}_{i,t}$ for all $i\in[m]$. Then, the parameter server calculates the mean of these $\tilde{g}_{i,t}$ and denotes it as $\tilde{g}_{t}=\frac{1}{m}\sum_{i=1}^{m}\tilde{g}_{i,t}$ (Line 11). Next (Line 13), the server updates $z_{0}$ in the dual space to $z_{t}=z_{0}-\sum_{k=1}^{t}\eta_{k}\tilde{g}_{k}.$ It then maps $z_{t}$ back to the point $\theta_{t+1}$ in the primal space (Line 15). The illustration of the primal and dual updates in RDGD is presented in Fig. 2. We let
$\nabla\psi^{*}(z_{t})=\operatorname*{arg\,min}_{u\in\Theta}\Big{\{}\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k},u-\theta_{k}\rangle+\ D_{\psi}(u,\theta_{0})\Big{\}}.$ (7)
This process can be considered as a modified version of lazy mirror descent [27]. The intuition behind RDGD  is that in (7), we are minimizing an approximation of the loss function, regularized by the Bregman divergence $D_{\psi}(u,\theta_{0})$. Introducing this regularizer prevents the updated point $\theta_{t+1}$ from straying too far from the current point $\theta_{t}$ when there are corruptions. Also, we denote $\hat{\theta}_{t}=\frac{H_{t-1}}{H_{t}}\hat{\theta}_{t-1}+\frac{\eta_{t}}{H_{t}}\theta_{t},$ (where $H_{t}=\sum_{k=1}^{t}\eta_{k}$) as the output at iteration $t$ (Line 14). We note that $\hat{\theta}_{t}$ is the sum of the weighted gradients based on the history of gradients. As the analysis shows, this ensures that the final output of RDGD $\hat{\theta}_{T}$ is robust against noisy updates.
III-B Convergence Analysis
We apply the approximate duality gap technique proposed in [28] to analyze the convergence of RDGD. We will show how to construct an upper bound $G_{t}$ of the suboptimality gap $L(\hat{\theta}_{t})-L({\theta^{*}})$, where $\hat{\theta}_{t}$ is the output of a first-order method at the $t$-th iteration. Denote $G_{t}=U_{t}-L_{t}$, where $U_{t}\geqslant L(\hat{\theta}_{t})$ is an upper bound on $L(\hat{\theta}_{t})$ and $L_{t}\leqslant L({\theta^{*}})$ is a lower bound on $L({\theta^{*}})$. Define $H_{t}\coloneq\sum_{k=1}^{t}\eta_{k}$. For the lower bound $L_{t}$, by the convexity of $L$, we have
$L(\theta^{*})\geqslant\frac{\sum_{k=1}^{t}\eta_{k}L(\theta_{k})+\sum_{k=1}^{t}\eta_{k}\langle\nabla L(\theta_{k}),\theta^{*}-\theta_{k}\rangle}{H_{t}}.$ (8)
To express the gradient $\nabla L(\theta_{k})$ in relation to the corrupted gradients $\{\tilde{g}_{i,k}:i\in[m]\}$, we write $\nabla L(\theta_{k})$ as $\nabla L(\theta_{k})=\tilde{g}_{k}+\nabla L(\theta_{k})-\tilde{g}_{k}$. Also, by adding and subtracting $\frac{1}{H_{t}}D_{\psi}(\theta^{*},\theta_{0})$ in (8), we have
$H_{t}L(\theta^{*})\geqslant\sum_{k=1}^{t}\eta_{k}L(\theta_{k})+\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k},\theta^{*}-\theta_{k}\rangle$
$-\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k}\rangle+D_{\psi}(\theta^{*},\theta_{0})-D_{\psi}(\theta^{*},\theta_{0}).$
Next, we can obtain our lower bound by replacing $\theta^{*}$ by a minimization over $\Theta$:
$H_{t}L(\theta^{*})\geqslant\sum_{k=1}^{t}\eta_{k}L(\theta_{k})-\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k}\rangle$
$-D_{\psi}(\theta^{*},\theta_{0})+\min_{u\in\Theta}\Big{\{}\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k},u-\theta_{k}\rangle+D_{\psi}(u,\theta_{0})\Big{\}}.$
We then define $H_{t}L_{t}$ as the lower bound above, i.e.,
$L_{t}$ :-$\coloneq\frac{1}{H_{t}}\Big{\{}\sum_{k=1}^{t}\eta_{k}L(\theta_{k})-\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k}\rangle:$
$\hskip 85.35826pt-D_{\psi}(\theta^{*},\theta_{0})+\min_{u\in\Theta}h_{t}(u)\Big{\}},$
where
$h_{t}(u)\coloneq\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k},u-\theta_{k}\rangle+D_{\psi}(u,\theta_{0}).$ (9)
The promised upper bound is chosen as $U_{t}=L(\hat{\theta}_{t})$, where $\hat{\theta}_{t}$ is a weighted average of the points $(\theta_{k})_{k=1}^{t}$ generated from RDGD  up to $t$-th iteration (specified in Line 14). Given the forms of $U_{t}$ and $L_{t}$, it is evident that $L(\hat{\theta}_{t})-L(\theta^{*})\leqslant G_{t}$. Therefore, demonstrating the convergence of the algorithm primarily is tantamount to establishing a bound for $G_{t}$. We need to track the change of $E_{t}=H_{t}G_{t}-H_{t-1}G_{t-1}$. To facilitate subsequent analyses, we express $G_{t}$ as:
$G_{t}=\frac{H_{1}}{H_{t}}G_{1}+\frac{\sum_{k=2}^{t}E_{k}}{H_{t}}.$
According to RDGD, $\theta_{t+1}=\operatorname*{arg\,min}_{u\in\Theta}h_{t}(u)$. For our analysis, we first bound $h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t})$ to track the change of the lower bound $L_{t}$ in the following lemma. The proof is provided in Appendix A.
Lemma 1
For all $t\geqslant 1$, the following holds:
$h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t})\geqslant\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle+D_{\psi}(\theta_{t+1},\theta_{t}).$
Now, we can derive an upper bound on $H_{t}G_{t}-H_{t-1}G_{t-1}$ based on Lemma 1. The proof is provided in Appendix B.
Lemma 2
For all $t\geqslant 1$, the following holds:
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle\\ -\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t}).$
Finally, we state the convergence rate of RDGD.
Theorem 1
Let $0<M<\infty$ be such that for all $(x,y)\in\mathbb{R}^{p}\times\mathbb{R}$, $\ L(x,y;\cdot)$ is $M$-smooth. Let $\theta_{0}\in\Theta$ be an arbitrary initial point. Let $\{(\theta_{t},z_{t},\hat{\theta}_{t})\}_{t=1}^{T}$ evolve according to RDGD (Algorithm 1) for some $\mu$-strongly convex function $\psi$, where $R_{\Theta}=\max_{\theta\in\Theta}\|\theta-\theta^{*}\|_{2}$. For a fixed number of iterations $T$ and stepsize $\eta_{k}=\frac{1}{\sqrt{T}}$, we have for all $k\geqslant 1$:
$\mathbb{E}[L(\hat{\theta}_{T})-L(\theta^{*})]$ $\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{T}}+\frac{M^{2}}{2\mu}\cdot\frac{1}{\sqrt{T}}$
$+\frac{R_{\Theta}C(T)}{m}\cdot\frac{1}{\sqrt{T}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
The first two terms $\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{T}}$ and $\frac{M^{2}}{2\mu}\cdot\frac{1}{\sqrt{T}}$ are similar to what one would expect for mirror descent [29]. Since we assume that $\Theta$ is bounded, $R_{\Theta}$ is finite. The effect of the adversarial corruption is manifested in the term $\frac{R_{\Theta}C(T)}{m}\cdot\frac{1}{\sqrt{T}}$, where recall that $C(T)$ (defined in (5)) represents the total corruption among all $m$ workers over $T$ iterations. Even though we do not have a lower bound on the gap $\mathbb{E}[L(\hat{\theta}_{T})-L(\theta^{*})]$, the scaling $O\big{(}\frac{C(T)}{m\sqrt{T}}\big{)}$ is intuitively order-optimal as, based on (5), if each worker suffers a constant amount of corruption at each time, then $C(T)=\Theta(m\sqrt{T})$. Besides, if the variances of the noises in both downlink and uplink channels are bounded across all iterations (i.e., $\sigma_{t}^{2}\leqslant\bar{\sigma}^{2}$ for all $t$ and some $\bar{\sigma}^{2}<\infty$), the effect of noisy channels on the bound is the constant $(M+1)\sqrt{p}\sigma R_{\Theta}$. We provide the proof of Theorem 1 in Appendix C.
If the number of iterations $T$ is not known a priori, then $\eta_{k}$ cannot depend on $T$. In this case, we can set $\eta_{k}=\frac{1}{\sqrt{k}}$ for all $k\geqslant 1$. A simple corollary of Theorem 1 says that under the same assumptions, for all $t\geqslant 1$,
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{t}}+\frac{M^{2}}{2\mu}\cdot\frac{\log t}{\sqrt{t}}$
$+\frac{R_{\Theta}C(t)}{m}\cdot\sqrt{\frac{\log t}{t}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
The only difference vis-à-vis the convergence rate specified in Theorem 1 is the corruption term $O\big{(}\sqrt{\frac{\log t}{t}}\big{)}$, which is slightly worse, by a factor of $\sqrt{\log t}$, than its counterpart in Theorem 1. This is due to our incognizance of the number of iterations $T$.
IV The Strongly Convex Case
In this section, we make a further assumption that the loss function $L$ is strongly convex (in addition to being smooth), and consequently, we are able to improve on the convergence rate in Theorem 1. Our extended algorithm is known as RDGD-SC (Algorithm 2), where the suffix SC stands for Strongly Convex. In essence, assuming that the loss function $L$ is $\alpha$-strongly convex, the only step that differs from Algorithm 1 is Line 15 therein, which now reads:
$\theta_{t+1}$ $=\operatorname*{arg\,min}_{u\in\Theta}\Big{\{}\sum_{k=1}^{t}\eta_{k}\left(\langle\tilde{g}_{t},u-\theta_{k}\rangle+\frac{\alpha}{2}\|u-\theta_{k}\|^{2}\right)$
$\hskip 56.9055pt+D_{\psi}(u,\theta_{0})\Big{\}}.$ (10)
The purpose of the quadratic term $\frac{\alpha}{2}\|u-\theta_{k}\|^{2}$ in (10) is to exploit the strong convexity of the loss function $L$ to ensure that the expected suboptimality gap decays faster; typically exponentially fast in $t$.
Algorithm 2 Distributed Gradient Descent Algorithm Robust to Corruptions over Noisy Channels for Strongly-Convex Functions (RDGD-SC)
1:Initialization: Parameter vector $\theta_{0}\in\Theta$, algorithm parameters $\eta_{t}$, dual vector $z_{0}=\nabla\psi(\theta_{0})$ with $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$, $\theta_{1}=\theta_{0}$ and $\hat{\theta}_{0}=\theta_{0}$.
2:for $t=1,2,\ldots$ do
3:   Parameter server: Send $\theta_{t}$ to all the workers over the downlink noisy channel.
4:   for all $i\in[m]$ do in parallel
5:      Worker $i$:
6:      Receive $\theta_{i,t}=\theta_{t}+v_{t}^{(i)}$ and calculate the corrupted local gradient $g^{\prime}_{i,t}$ in (2).
7:      Send $\bar{g}_{i,t}=g^{\prime}_{i,t}+\varepsilon_{i,t}$ to parameter server via the noisy uplink channel.
8:   end for
9:   Parameter server:
10:   Receive $\tilde{g}_{i,t}=\bar{g}_{i,t}+w_{t}^{(i)}$, for all $i\in[m]$.
11:   Compute mean gradient $\tilde{g}_{t}\leftarrow\frac{1}{m}\sum_{i=1}^{m}\tilde{g}_{i,t}$.
12:   Update model parameter:
13:   $z_{t}=z_{t-1}-\eta_{t}\tilde{g}_{t}$, i.e., $z_{t}=-\sum_{k=1}^{t}\eta_{k}\tilde{g}_{k}+z_{0}$.
14:   $\hat{\theta}_{t}=\frac{H_{t-1}}{H_{t}}\hat{\theta}_{t-1}+\frac{\eta_{t}}{H_{t}}\theta_{t}$, where $H_{t}=\sum_{k=1}^{t}\eta_{k}$.
15:   $\theta_{t+1}=\operatorname*{arg\,min}_{u\in\Theta}\{\sum_{k=1}^{t}\eta_{k}(\langle\tilde{g}_{t},u-\theta_{k}\rangle+\frac{\alpha}{2}\|u-\theta_{k}\|^{2})+D_{\psi}(u,\theta_{0})\}$.
16:end for
17:Output: The sequence of parameter vectors $\{\hat{\theta}_{t}\}_{t=1}^{\infty}$.
To analyze the convergence behavior of RDGD-SC, we also apply the approximate duality gap technique used in Section III. For the sake of brevity, we only emphasize the parts that are different from Section III in this section. For the lower bound $L_{t}$, exploiting strongly convexity instead of regular convexity, we redefine it as
$L_{t}\coloneq\frac{1}{H_{t}}\Big{\{}\sum_{k=1}^{t}\eta_{k}L(\theta_{k})$ $-\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k}\rangle$
$-D_{\psi}(\theta^{*},\theta_{0})+\min_{u\in\Theta}h_{t}(u)\Big{\}}$
where $h_{t}$ is also redefined as
$h_{t}(u)\coloneq\sum_{k=1}^{t}\eta_{k}\Big{(}\langle\tilde{g}_{k},u-\theta_{k}\rangle+\frac{\alpha}{2}\|u-\theta_{k}\|_{2}^{2}\Big{)}+D_{\psi}(u,\theta_{0}).$ (11)
The template for our analysis is similar to that in the smooth case. The following definitions are identical, starting with $\theta_{t+1}=\mathrm{argmin}_{u\in\Theta}h_{t}(u)$ and $G_{t}=U_{t}-L_{t}$. Additionally, we need to track the change of $E_{t}=H_{t}G_{t}-H_{t-1}G_{t-1}$, which leads to $G_{t}=\frac{H_{1}}{H_{t}}G_{1}+\frac{\sum_{k=2}^{t}E_{k}}{H_{t}}$. The following lemma bounds $H_{1}G_{1}$ and its proof is provided in Appendix D.
Lemma 3
Let $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$, the following holds:
$H_{1}G_{1}\leqslant\frac{\eta_{1}\alpha}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle+\frac{\eta_{1}M^{2}}{4\alpha}.$
The following lemma bounds $h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t})$ and serves to track the change of the lower bound $L_{t}$. Its proof is provided in Appendix E.
Lemma 4
Let $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$. For all $t\geqslant 1$,
$h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t})\geqslant\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle\\ +\frac{(H_{t}+\eta_{t})\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$
Finally, the following lemma whose proof is provided in Appendix F provides a bound of $H_{t}G_{t}-H_{t-1}G_{t-1}$.
Lemma 5
Let $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$ and $0\leqslant\frac{\eta_{t}}{H_{t}}\leqslant\frac{\alpha}{M}$. Then, for all $t\geqslant 1$,
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle+\frac{\eta_{t}M^{2}}{2\alpha}.$
Equipped with these preliminary results, the convergence rate of RDGD-SC can be quantified as follows.
Theorem 2
Let $0<\alpha\leqslant M<\infty$ be such that for all $(x,y)\in\mathbb{R}^{p}\times\mathbb{R}$, $\ L(x,y;\cdot)$ is $M$-smooth and $\alpha$-strongly convex. let $\theta_{0}\in\Theta$ be an arbitrary initial point. Let $\{(\theta_{t},z_{t},\hat{\theta}_{t})\}_{t=1}^{T}$ evolve according to RDGD-SC (Algorithm 2) for $\psi(x)=\frac{\eta_{1}\alpha}{2}\|x\|_{2}^{2}$. Set $R_{\Theta}=\max_{\theta\in\Theta}\|\theta-\theta^{*}\|_{2}$. Let $\eta_{1}=1$ and $\eta_{k}$ satisfy $\eta_{k}\leqslant\frac{\alpha}{M}H_{k}$ for all $k\geqslant 2$. Then for all $t\geqslant 2$:
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant\Big{(}\prod_{k=2}^{t}\big{(}1-\frac{\eta_{k}}{H_{k}}\big{)}\Big{)}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}$
$+\frac{R_{\Theta}C(t)}{m}\cdot\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$ (12)
Proof:
We write $G_{t}=\frac{H_{1}}{H_{t}}G_{1}+\frac{\sum_{k=2}^{t}E_{k}}{H_{t}}$ as
$G_{t}$ $=\frac{H_{1}}{H_{2}}\frac{H_{2}}{H_{3}}\dotsm\frac{H_{t-1}}{H_{t}}G_{1}+\frac{\sum_{k=2}^{t}E_{k}}{H_{t}}$
$=\Big{(}1-\frac{\eta_{2}}{H_{2}}\Big{)}\Big{(}1-\frac{\eta_{3}}{H_{3}}\Big{)}\ldots\Big{(}1-\frac{\eta_{t}}{H_{t}}\Big{)}G_{1}+\frac{\sum_{k=2}^{t}E_{k}}{H_{t}}.$
The remaining parts of the proof are essentially the same as those for Theorem 1. ∎
Observe that we have the flexibility of setting different stepsize sequences $\{\eta_{k}\}_{k=1}^{\infty}$ to achieve different convergence rates in (12). Motivated by this observation, we derive the following corollary of Theorem 2. The proof is presented in Appendix G.
Corollary 1
(of Theorem 2). For $k\geqslant 2$, setting
  • $\eta_{k}=\frac{\alpha}{M}H_{k}$, we have
    $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}$
    $+\frac{R_{\Theta}C(t)}{m}+(M\!+\!1)\sqrt{p}\sigma R_{\Theta}.$ (13)
  • $\eta_{k}=\frac{2}{k+1}H_{k}$, we have
    $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ $\leqslant O\Big{(}\frac{\|\theta^{*}-\theta_{0}\|_{2}^{2}}{t^{2}}+\frac{1}{\sqrt{t}}\cdot\frac{R_{\Theta}C(t)}{m}\Big{)}$
    $\quad+\frac{M^{2}}{2\alpha}+(M+1)\sqrt{p}\sigma R_{\Theta}.$ (14)
If we set $\eta_{k}=\frac{\alpha}{M}H_{k}$ (the first case in Corollary 1), we notice that the leading term $\left(1-\frac{\alpha}{M}\right)^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}$ resembles the convergence behavior typically observed in the analysis of smooth and strongly convex functions, in scenarios without corruption and noise [30]. This term decays exponentially fast as $t$ grows. However, there is an additional term of the form $\frac{R_{\Theta}C(t)}{m}$ that represents the effect of the corruption and does not vanish with $t$. Alternatively, if we set $\eta_{k}=\frac{2}{k+1}H_{k}$ (the second case of Corollary 1), the leading term vanishes at a slower rate of $t^{-2}$ but the corruption term vanishes at a rate $\frac{C(t)}{m\sqrt{t}}$. This naturally begs the question of whether we can design a “best of both worlds” algorithm to exploit the benefits of both settings of the stepsize $\eta_{k}$ to accelerate the overall convergence.
V Corruption Reduction by Exploiting the Best of Both Worlds
Based on the results from Section IV, we observe that for certain natural choices of the stepsize such as $\eta_{k}=\frac{\alpha}{M}H_{k}$ with $H_{k}=\sum_{\ell=1}^{k}\eta_{\ell}$, the adversarial corruption results in the upper bound on the suboptimality gap in (13) stagnating at a positive, non-vanishing constant. This is clearly undesirable. In this section, we explore how to mitigate this corruption accumulation by proposing an algorithm that switches the stepsize schedule at an appropriate transition time. The main idea is to first exploit the exponential decrease of the expected suboptimality gap by setting $\eta_{k}=\frac{\alpha}{M}H_{k}$ until the corruption accumulation begins to dominate the convergence behavior. At this point, we switch to a different stepsize schedule (i.e., $\eta_{k}=\frac{2}{k+1}H_{k}$), allowing the corruption to be gradually amortized and thus enabling the hybrid algorithm to further reduce the suboptimality gap as $t$ grows.
In this section, for clarity, we specify the total corruption budget as $C(t)=mt^{r}$ for some fixed $r\in(0,1/2)$. We observe from (13) that the suboptimality gap initially decreases and then increases due to the accumulation of the corruption $C(t)$ when we set $\eta_{k}=\frac{\alpha}{M}H_{k}$. Hence, we determine the optimal transition time $t_{0}$ corresponding to the point where the minimum upper bound of suboptimality gap is achieved. The proof of the following lemma is provided in Appendix H.
Lemma 6
If $\eta_{k}=\frac{\alpha}{M}H_{k}$, the minimum of the upper bound of $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ in (13) with respect to $t\in\mathbb{N}$ is attained at
$t_{0}=\Big{\lceil}-\frac{(1-r)M}{\alpha}W_{-1}\big{(}B(r)\big{)}\Big{\rceil},$ (15)
where $B(r):=-\frac{\alpha}{(1-r)M}\big{(}\frac{2Mr}{\alpha^{2}R_{\Theta}\exp(\alpha/M)}\big{)}^{\frac{1}{1-r}}$, and $W_{-1}$ denotes the Lambert W function, i.e., the inverse of $f(w)=we^{w}$.
Note from Lemma 6 that the transition time $t_{0}$ can be analytical determined based on known parameters of the problem such as $M$, $\alpha$, and $R_{\Theta}$. Inspired by Lemma 6, we propose the Algorithm 3 that is based on RDGD but with a novel restart mechanism (RDGD-Restart). RDGD-Restart exploits the exponential decrease in the suboptimality gap until the effect of the accumulation of the corruption starts to dominate the overall bound in (13). To mitigate this, we then switch to a stepsize sequence that decays more slowly to prevent excessive corruption accumulation. This then leads to the bound in (14) dominating the overall convergence behavior subsequently.
Algorithm 3 Distributed Gradient Descent Algorithm Robust to Corruptions over Noisy Channels with Restart (RDGD-Restart)
1:Initialization: Parameter vector $\theta_{0}\in\Theta$, $\eta_{1}=1$, dual vector $z_{0}=\nabla\psi(\theta_{0})$ with $\psi(x)=\frac{\alpha}{2}\|x\|_{2}^{2}$, $\theta_{1}=\theta_{0}$ and $\hat{\theta}_{0}=\theta_{0}$ (defined in (15)).
2:Stage 1: Run RDGD-SC (Algorithm 2) with $\eta_{k}=\frac{\alpha}{M}H_{k}$ from $k\geqslant 2$. Terminate when $k=t_{0}$.
3:Stage 2: Restart RDGD-SC taking $\theta_{t_{0}}$ as the initial point and change the stepsize to $\eta_{k}=\frac{2}{k+1}H_{k}$ for all $k\geqslant t_{0}+1$.
4:Output: The parameter vector $\hat{\theta}_{t}$.
We now bound the suboptimality gap $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ of RDGD-Restart. The proof is presented in Appendix I.
Theorem 3
Under the same setting as Theorem 2 and when corruption budget is paramterized as $C(t)=mt^{r}$ for $r\in(0,1/2)$, for all $t>t_{0}$ (defined in (15)), the convergence rate of RDGD-Restart in Algorithm 3 is:
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\frac{\alpha}{2}\Big{(}1-\frac{\alpha}{M}\Big{)}^{t_{0}-1}\frac{t_{0}(t_{0}+1)}{t(t+1)}\|\theta^{*}-\theta_{0}\|_{2}^{2}\\ +R_{\Theta}t^{r}\cdot A(t,t_{0})+\frac{M^{2}}{2\alpha}+(M+1)\sqrt{p}\sigma R_{\Theta},$
where $A(t,t_{0})\coloneq\sqrt{\left(\frac{t_{0}(t_{0}+1)}{t(t+1)}\right)^{2}\!+\!\left(\frac{2}{t(t+1)}\right)^{2}\sum_{k=t_{0}+1}^{t}k^{2}}$.
As $t\to\infty$, $A(t,t_{0})$ decays as $O(1/\sqrt{t})$. We observe from the first term of the bound that in the initial rounds (when $t\leqslant t_{0}$), the suboptimality gap decays rapidly due to the exponentially decaying term with contraction factor $1-\frac{\alpha}{M}$. When $t>t_{0}$, we observe from the second term that the effect of the corruption is amortized because $r<1/2$ and $A(t,t_{0})=O(1/\sqrt{t})$.
VI Numerical Experiments
In this section, we illustrate the performance of RDGD and its variants through several numerical experiments, including least squares regression and support vector machine (SVM) classification on synthetically generated datasets, and softmax classification on the MNIST handwritten digits dataset [31]. We compare our proposed algorithms with the vanilla distributed gradient descent (DGD) algorithm, i.e., the parameter vector is simply updated as $\theta_{t+1}=\theta_{t}-\eta_{t}\tilde{g}_{t}$ at the $t$-th iteration. Note that DGD does not attempt to correct for the presence of corruptions; hence, we expect that it is not robust.
VI-A Least Squares Regression on Synthetic Datasets
We start with least squares regression on synthetic datasets. The datasets contain $N=10,000$ data samples. In particular, the loss function of the least-squares regression is as $L(x,y;\theta)=\frac{1}{2}(x^{\top}\theta-y)^{2}$ where $\theta\in\mathbb{R}^{p}$ is the parameter vector with dimensionality $p=20$. Then, the loss function computed over the entire dataset $\mathcal{Z}$ is $L(\theta)=\frac{1}{2N}\sum_{i=1}^{N}(x_{i}^{\top}\theta-y_{i})^{2}$. We consider one parameter server and $m=20$ workers.
Each element of the feature vectors $x_{i}$ is drawn from the standard Gaussian distribution $\mathcal{N}(0,1)$. We also randomly generate each dimension of the unknown parameter $\theta^{*}$ from $\mathcal{N}(0,1)$. The scalar responses $y_{i}$ are obtained by taking the inner product of the feature vectors $x_{i}$ and the unknown parameter $\theta^{*}$, followed by adding standard Gaussian noise $\mathcal{N}(0,1)$. These data samples are then uniformly distributed to all $20$ workers. We set the variances of the Gaussian noise in both downlink and uplink channels to $\sigma^{2}=0.5$.
The performance metric is the suboptimality gap $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$. The results reported are averaged over 100 independent trials. We also report the standard deviations across these 100 experiments. The bound of the total corruption until time $t$ is set to $C(t)=20t^{0.4}$. As it is difficult to find the worst case corruption pattern, we use the following procedure as a proxy to generate the corruption $\varepsilon_{i,t}$ at each worker $i\in[m]$ and at each time $t$. In particular, let $\Gamma(\tau)=\sum_{k=1}^{\tau}c(k)^{2}$ denote the corruption budget used up to time and including $n$. Then, the available corruption budget at time $t$ is $\sqrt{C(t)^{2}-\Gamma(t-1)}$. Then, we distribute this available corruption budget to each worker randomly. The sign of adversarial corruption at each worker $\operatorname{sign}(\varepsilon_{i,t})$ is set to $-\operatorname{sign}(\tilde{g}_{i,t})$ for all $i\in[m]$. Also, the stepsize is set to $\eta_{t}=t^{-2}$.
Figure 3: Performances of RDGD  and DGD  for least squares regression
Let $X\in\mathbb{R}^{N\times p}$ be the matrix where each row contains the feature vector $x_{i}^{\top}$. Define $\lambda_{\mathrm{max}}(Q)$ and $\lambda_{\mathrm{min}}(Q)$ as the maximum and minimum eigenvalues of the matrix $Q$ respectively. We run RDGD (Algorithm 1) with the function $\psi(x)=\frac{M}{2}\|x\|_{2}^{2}$, where $M=\lambda_{\mathrm{max}}(\frac{1}{N}X^{\top}X)$. As shown in Fig. 3, our proposed algorithm achieves a significantly smaller suboptimality gap more quickly and consistently compared to the traditional DGD algorithm.
Next, we verify the efficacy of the restart mechanism RDGD-Restart  (in Algorithm 3) when the loss function is smooth and strongly convex. In particular, we modify the loss function of least-squares regression by adding an $\ell_{2}$-squared regularization term, i.e., $L(x,y;\theta)=\frac{1}{2}(x^{\top}\theta-y)^{2}+\frac{\lambda}{2}\|\theta\|_{2}^{2}$, where we set $\lambda=0.01$. Now, we run RDGD-Restart for this strongly convex loss function. We set the strongly convex parameter as $\alpha=\lambda_{\mathrm{min}}(\frac{1}{N}X^{\top}X+\lambda\mathbf{I}_{p})$ and the smooth parameter as $M=\lambda_{\mathrm{max}}(\frac{1}{N}X^{\top}X+\lambda\mathbf{I}_{p})$ respectively. The variance of the Gaussian noises in the uplink and downlink is set to $\sigma^{2}=0.1$. The function $\psi(x)=\frac{\alpha}{2}\|x\|_{2}^{2}$.
Figure 4: Comparison of the performances of RDGD-SC with different stepsizes $\{\eta_{k}\}$ to RDGD-Restart
As shown in Fig. 4, if one chooses the stepsize sequence $\eta_{k}=\frac{\alpha}{M}H_{k}$, the adversarial corruption accumulates after a certain point, resulting in a stagnation of the suboptimality gap. However, it is also worth noting that the gap decreases much more rapidly when $\eta_{k}=\frac{\alpha}{M}H_{k}$ compared to when $\eta_{k}=\frac{2}{k+1}H_{k}$ initially. In our proposed algorithm with a restart mechanism RDGD-Restart, the gap decreases rapidly until a transition time $t_{0}$ which can be determined analytically. After the time $t_{0}+1$, we switch the stepsize sequence from $\eta_{k}=\frac{\alpha}{M}H_{k}$ to $\eta_{k}=\frac{2}{k+1}H_{k}$. As can be seen, the gap subsequently decreases with $t$. Our proposed algorithm RDGD-Restart converges faster than the $\eta_{k}=\frac{2}{k+1}H_{k}$ case throughout the entire process, indicating a smaller time complexity to achieve the same suboptimality gap. The experimental results presented in Fig. 4 thus corroborate our theoretical analysis in Section V. Hence, our proposed algorithm with a restart mechanism RDGD-Restart  successfully takes advantage of the exponential decrease in the gap until corruption accumulation occurs, then mitigates the adverse effects of the accumulation by switching to another stepsize sequence.
VI-B Classification using L2 Support Vector Machines
Next we verify the efficacy of RDGD  on a binary classification task with a loss function that is different from that used in the previous section. Specifically, we consider L2-SVM [^2] We consider L2-SVM instead of the usual L1-SVM since the loss function is smooth for the former. [32, 33] and we use the accuracy of the binary classification task as our performance metric. The label $y_{i}$ is generated such that it takes values from the set $\{1,-1\}$ with equal probability. For each $y_{i}$, the corresponding feature vector $x_{i}$ is generated from a multivariate normal distribution. If $y_{i}=1$, $x_{i}$ is drawn from $\mathcal{N}(\bm{\mu},\gamma\mathbf{I}_{p})$, and if $y_{i}=-1$, $x_{i}$ is drawn from $\mathcal{N}(-\bm{\mu},\gamma\mathbf{I}_{p})$, where $\bm{\mu}$ is the vector of all ones of length $p$ and $\gamma=4$.
We generated $N=10,000$ data samples which are then split into $8,000$ samples for training and $2,000$ samples for testing. The primal objective function used in L2-SVM is $L(x,y;\theta)=\max\{0,1-y(x^{\top}\theta)\}^{2}+\frac{\lambda}{2}\|\theta\|_{2}^{2}$, where $\lambda$ is set to $0.1$. We compare RDGD with vanilla DGD under varying levels of corruption budgets $C(T)$, with a fixed noise variance $\sigma^{2}=1$ in the (uplink and downlink) communication channels. Specifically, the total corruption bound up to time $T$ is set as $C(T)=20T^{r}$, where $r\in\{0.25,0.3,0.35\}$. The total corruption budget is uniformly allocated to each iteration and worker, i.e., $c(t)=C(T)/\sqrt{T}$, for all $t\in[T]$.
Figure 5: RDGD for L2-SVM classification
As shown in Fig. 5, the proposed RDGD converges rapidly and achieves high accuracy at any level of the corruption budget. However, the accuracy of the classification of the vanilla DGD algorithm is lower than RDGD and is highly sensitive to the level of the corruption budget. In particular, the performance of DGD fails when the corruption budget exceeds a certain level (e.g., $r=0.35$). In Fig. 6, we display the decision boundaries for the first two dimensions of the parameter vectors $\hat{\theta}_{\rm{RDGD}}$ and $\hat{\theta}_{\rm{DGD}}$ obtained by RDGD and DGD respectively when $r=0.35$. Note that $\hat{\theta}_{\rm{RDGD}}$ is almost identical to $\theta^{*}$. However, $\hat{\theta}_{\rm{DGD}}$ deviates significantly from $\theta^{*}$, showing that DGD is far from robust.
Figure 6: Decision Boundaries for L2-SVM classification
VI-C Softmax Classification on the MNIST Dataset
In our final set of experiments, we consider a multi-class classification task on the classical MNIST dataset, which contains 10 classes (i.e., the number of values that each label $y_{i}$ can take on is 10), 60,000 training samples and 10,000 test samples. The performance metric is the classification accuracy on the test samples. The dimensionality $p$ of the data points is $784=28\times 28$ and the number of workers $m$ is set to $20$. We use the cross-entropy loss to optimize the softmax classifier.
Figure 7: RDGD for the MNIST dataset with different levels of corruption
In Fig. 7, we compare RDGD with vanilla DGD under varying levels of corruption budgets $C(T)$, with a fixed noise variance $\sigma^{2}=0.5$ in the communication channels. Specifically, the total corruption bound up to time $T$ is set as $C(T)=100T^{r}$, where $r\in\{0.2,0.3,0.4\}$. The total corruption budget is also uniformly allocated to each iteration and each worker as in the previous subsection. As shown in Fig. 7, RDGD  outperforms the traditional DGD algorithm across all levels of the corruption budget. Furthermore, we observe that RDGD achieves higher accuracy as the value of $r$ and hence, the amount of corruption, decreases. The vanilla DGD algorithm reduces to random guessing (success rate $\approx 10\%$) when the corruption budget exceeds a certain level (i.e., $r=0.4$).
VII Conclusion
In this paper, we considered a robustification of the standard distributed gradient descent algorithm used extensively in the parallelization of massive datasets in the modern era. In addition to standard channel noises on the uplink and downlink of the communication channels, we also considered mitigating the effects of adversarial corruptions at the workers via a mirror descent-type algorithm RDGD and its strongly convex extension RDGD-SC. Different convergence rates of the suboptimality gap were obtained by setting different stepsizes in RDGD-SC. We also provided a theoretical analysis of a unified algorithm RDGD-Restart that attains the “best-of-both-worlds” with regard to the choice of the different stepsizes in RDGD-SC. Extensive numerical experiments show that (i) RDGD  performs much better, in terms of the suboptimality gap and classification accuracy, than standard DGD in reducing the effect of the corruption and (ii) the restarting mechanism in RDGD-Restart  is effective in speeding up the overall convergence rate for strongly convex loss functions.
We may consider the following as avenues for future research. First, to model the scenario in which some, but not all, workers are compromised, we may redefine $c_{t}$ in (4) as $c_{t}=\sum_{i=1}^{m}\|\varepsilon_{i,t}\|_{2}$. This encourages sparsity among the $\varepsilon_{i,t}$’s for fixed $t$. Second, we may derive information-theoretic lower bounds on the suboptimality gap in terms of the amount of corruption $C(T)$ to ascertain the tightness of the upper bounds.
Appendix A Proof of Lemma 1
Proof:
By the definition of $h_{t}$ in (9), we have $h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t+1})\geqslant\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle$.
Then, based on the definition of the Bregman divergence, we have $h_{t-1}(\theta_{t+1})-h_{t-1}(\theta_{t})=\langle\nabla h_{t-1}(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+D_{h_{t-1}}(\theta_{t+1},\theta_{t})$ and $D_{h_{t-1}}(\theta_{t+1},\theta_{t})=D_{\psi}(\theta_{t+1},\theta_{t})$. However, since $\theta_{t}=\mathrm{argmin}_{u\in\Theta}h_{t-1}(u)$, we have $\langle\nabla h_{t-1}(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\geqslant 0$. Thus,
$h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t})\geqslant\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle+D_{\psi}(\theta_{t+1},\theta_{t})$
as desired. ∎
Appendix B Proof of Lemma 2
Proof:
Based on Lemma 1, we have:
$H_{t}L_{t}-H_{t-1}L_{t-1}$ $\geqslant\eta_{t}L(\theta_{t})-\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t}\rangle$
$\qquad+\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle+D_{\psi}(\theta_{t+1},\theta_{t})$
$=\eta_{t}L(\theta_{t})-\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle$
$\;+\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+D_{\psi}(\theta_{t+1},\theta_{t}).$
Then, it follows that:
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})-\eta_{t}L(\theta_{t})\\ +\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle-\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\\ -D_{\psi}(\theta_{t+1},\theta_{t}).$
From RDGD in Algorithm 1 and the convexity of the loss function $L$, we can write $H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})-\eta_{t}L(\theta_{t})$ as:
$\eta_{t}L(\hat{\theta}_{t})+H_{t-1}(L(\hat{\theta}_{t})-L(\hat{\theta}_{t-1}))-\eta_{t}L(\theta_{t})$
$\leqslant\eta_{t}(L(\hat{\theta}_{t})-L(\theta_{t}))+H_{t-1}\langle\nabla L(\hat{\theta}_{t}),\hat{\theta}_{t}-\hat{\theta}_{t-1}\rangle$
$=\eta_{t}(L(\hat{\theta}_{t})-L(\theta_{t}))-\eta_{t}\langle\nabla L(\hat{\theta}_{t}),\hat{\theta}_{t}-\theta_{t}\rangle$
$\leqslant\eta_{t}\langle\nabla L(\hat{\theta}_{t}),\hat{\theta}_{t}-\theta_{t}\rangle-\eta_{t}\langle\nabla L(\hat{\theta}_{t}),\hat{\theta}_{t}-\theta_{t}\rangle$
$=0.$ (16)
Hence, we conclude that
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle\\ -\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t})$
as desired. ∎
Appendix C Proof of Theorem 1
Proof:
Based on Lemma 2, we have
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant\eta_{t}\langle\tilde{g}_{t}-L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle\\ -\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t}).$
By the definition of the Bregman divergence and the $\mu$-strong convexity of the function $\psi$, we have $D_{\psi}(\theta_{t+1},\theta_{t})\geqslant\frac{\mu}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$. Hence,
$-\eta_{t}\langle L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-D_{\psi}(\theta_{t+1},\theta_{t})\\ \leqslant\eta_{t}M\|\theta_{t+1}-\theta_{t}\|_{2}-\frac{\mu}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}\leqslant\frac{\eta_{t}^{2}M^{2}}{2\mu},$
where the second inequality follows from the fact that $2ab-b^{2}\leqslant a^{2}$.
Recall that $g_{i,k}=\frac{1}{|\mathcal{Z}_{i}|}\sum_{(x,y)\in\mathcal{Z}_{i}}\nabla L(x,y;\theta_{k})$. Hence, we can write $\nabla L(\theta_{k})=\frac{1}{m}\sum_{i=1}^{m}g_{i,k}$ and
$\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle$
$\!=\!\bigg{\langle}\frac{1}{m}\Big{(}\!\sum_{i=1}^{m}g^{\prime}_{i,t}\!+\!\sum_{i=1}^{m}\varepsilon_{i,t}\!+\!\sum_{i=1}^{m}w_{k}^{(i)}\Big{)}$
$\hskip 56.9055pt-\frac{1}{m}\!\sum_{i=1}^{m}g_{i,t},\theta^{*}\!-\!\theta_{k+1}\bigg{\rangle}$
$=\frac{1}{m}\bigg{\langle}\!\sum_{i=1}^{m}\!g^{\prime}_{i,t}\!-\!\sum_{i=1}^{m}\!g_{i,t},\theta^{*}\!-\!\theta_{k+1}\!\bigg{\rangle}\!+\!\frac{1}{m}\bigg{\langle}\!\sum_{i=1}^{m}\!\varepsilon_{i,t},\theta^{*}\!-\!\theta_{k+1}\!\bigg{\rangle}$
$\hskip 56.9055pt-\frac{1}{m}\bigg{\langle}\sum_{i=1}^{m}w_{k}^{(i)},\theta^{*}-\theta_{k+1}\bigg{\rangle}.$
Recall that $R_{\Theta}=\max_{\theta\in\Theta}\|\theta-\theta^{*}\|_{2}$. It follows that
$\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle$
$\leqslant MR_{\Theta}\|v_{k}^{(i)}\|_{2}+\frac{c_{k}R_{\Theta}}{m}+R_{\Theta}\|w_{k}^{(i)}\|_{2}.$ (17)
Then, we have
$\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle\\ \leqslant\frac{R_{\Theta}}{m}\sum_{k=1}^{t}\eta_{k}c_{k}+MR_{\Theta}\sum_{k=1}^{t}\eta_{k}\|v_{k}^{(i)}\|_{2}+R_{\Theta}\sum_{k=1}^{t}\eta_{k}\|w_{k}^{(i)}\|_{2}.$ (18)
Taking the expectation on both sides of (18) and assuming $\sigma^{2}_{k}\leqslant\sigma^{2}$ for all $k\geqslant 1$, we obtain
$\mathbb{E}\bigg{[}\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle\bigg{]}$
$\leqslant\frac{R_{\Theta}}{m}\langle\bm{\eta}_{t},\mathbf{c}_{t}\rangle\!+\!(M+1)\sqrt{p}\sigma R_{\Theta}\sum_{k=1}^{t}\eta_{k},$ (19)
where $\bm{\eta}_{t}=[\eta_{1},\ldots,\eta_{t}]^{\top}$ and $\mathbf{c}_{t}=[c_{1},\ldots,c_{t}]^{\top}$.
Since the overall bound on the total corruption up to and including time $t$ is $C(t)$ such that $\sqrt{\sum_{k=1}^{t}c_{k}^{2}}\leqslant C(t)$, by the Cauchy–Schwarz inequality, we have
$\mathbb{E}\bigg{[}\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle\bigg{]}$
$\leqslant\frac{R_{\Theta}C(t)}{m}\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}+(M+1)\sqrt{p}\sigma R_{\Theta}H_{t}.$ (20)
Next, we bound $H_{1}G_{1}$ as follows:
$H_{1}G_{1}=H_{1}U_{1}-H_{1}L_{1}$
$\leqslant-\eta_{1}\langle L(\theta_{1}),\theta_{2}-\theta_{1}\rangle-D_{\psi}(\theta_{2},\theta_{0})+D_{\psi}(\theta^{*},\theta_{0})$
$\qquad\qquad\qquad+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle$
$\leqslant\frac{\eta_{1}^{2}M^{2}}{2\mu}+D_{\psi}(\theta^{*},\theta_{0})+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle.$
Then, we conclude that
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{H_{t}}+\frac{M^{2}}{2\mu}\cdot\frac{\sum_{k=1}^{t}\eta_{k}^{2}}{H_{t}}\\ +\frac{R_{\Theta}C(t)}{m}\cdot\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
Recall that $T$ is the fixed and known number of iterations. Let $\eta_{k}=\frac{1}{\sqrt{T}}$ for all $k\geqslant 1$. Finally we obtain the desired bound as follows:
$\mathbb{E}[L(\hat{\theta}_{T})-L(\theta^{*})]\leqslant\frac{D_{\psi}(\theta^{*},\theta_{0})}{\sqrt{T}}+\frac{M^{2}}{2\mu}\cdot\frac{1}{\sqrt{T}}\\ +\frac{R_{\Theta}C(T)}{m}\cdot\frac{1}{\sqrt{T}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
Appendix D Proof of Lemma 3
Proof:
The initial lower bound $L_{1}$ multiplied by $H_{1}$ can be expressed as
$H_{1}L_{1}=\eta_{1}L(\theta_{1})+\eta_{1}\langle\tilde{g}_{1},\theta_{2}-\theta_{1}\rangle+\Big{(}\frac{\eta_{1}\alpha}{2}+\frac{\alpha_{0}}{2}\Big{)}\|\theta_{2}-\theta_{1}\|_{2}^{2}\\ -\frac{\mu_{0}}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}-\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{1}\rangle.$
The corresponding upper bound $U_{1}$ multiplied by $H_{1}$ can be written as
$H_{1}U_{1}=\eta_{1}L(\hat{\theta}_{1})=\eta_{1}L(\theta_{1}).$
Then, we have
$H_{1}U_{1}-H_{1}L_{1}=H_{1}G_{1}$
$\leqslant\frac{\alpha_{0}}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle$
$\qquad\quad-\eta_{1}\langle\nabla L(\theta_{1}),\theta_{2}-\theta_{1}\rangle-\Big{(}\frac{\eta_{1}\alpha}{2}+\frac{\alpha_{0}}{2}\Big{)}\Big{\|}\theta_{2}-\theta_{1}\|_{2}^{2}$
$\leqslant\frac{\alpha_{0}}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle+\frac{\eta_{1}^{2}M^{2}}{2(\eta_{1}\alpha\!+\!\alpha_{0})}$
$=\frac{\eta_{1}\alpha}{2}\|\theta^{*}-\theta_{1}\|_{2}^{2}+\eta_{1}\langle\tilde{g}_{1}-\nabla L(\theta_{1}),\theta^{*}-\theta_{2}\rangle+\frac{\eta_{1}M^{2}}{4\alpha}$
as desired. ∎
Appendix E Proof of Lemma 4
Proof:
By the definition of $h_{t}$ in (11), we have $h_{t}(\theta_{t+1})-h_{t-1}(\theta_{t+1})\geqslant\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle+\frac{\eta_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$. Then, based on the definition of the Bregman divergence, we have
$h_{t-1}(\theta_{t+1})-h_{t-1}(\theta_{t})=\langle\nabla h_{t-1}(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\\ +D_{h_{t-1}}(\theta_{t+1},\theta_{t}).$
Since $\theta_{t}=\mathrm{argmin}_{u\in\Theta}h_{t-1}(u)$, we have $\langle\nabla h_{t-1}(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\geqslant 0$. Then:
$h_{t-1}(\theta_{t+1})-h_{t-1}(\theta_{t})$
$\geqslant D_{h_{t-1}}(\theta_{t+1},\theta_{t})$
$=\frac{H_{t-1}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}+\frac{\eta_{1}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$=\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
as desired. ∎
Appendix F Proof of Lemma 5
Proof:
From Lemma 4, we have
$H_{t}L_{t}-H_{t-1}L_{t-1}$
$\geqslant\eta_{t}L(\theta_{t})+\eta_{t}\langle\tilde{g}_{t},\theta_{t+1}-\theta_{t}\rangle+\frac{(H_{t}+\eta_{t})\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\qquad\qquad\qquad-\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t}\rangle$
$=\eta_{t}L(\theta_{t})\!+\!\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\!+\!\frac{(H_{t}\!+\!\eta_{t})\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\qquad\qquad-\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle.$
Then, we have
$H_{t}G_{t}-H_{t-1}G_{t-1}\leqslant H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})\\ -\eta_{t}L(\theta_{t})-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-\!\frac{(H_{t}\!+\!\eta_{t})\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}\\ +\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle.$ (21)
Combining the fact that $\hat{\theta}_{t}=\frac{H_{t-1}}{H_{t}}\hat{\theta}_{t-1}+\frac{\eta_{t}}{H_{t}}\theta_{t}$ and (B), we have $H_{t}L(\hat{\theta}_{t})-H_{t-1}L(\hat{\theta}_{t-1})-\eta_{t}L(\theta_{t})\leqslant 0$. Hence, we may rewrite the right-hand side of (21) as follows
$\eta_{t}(L(\theta_{t+1})-L(\theta_{t}))-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle\\ -\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}+\eta_{t}(L(\theta_{t})-L(\theta_{t+1}))-\!\frac{\eta_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}\\ +\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle.$
Combining the fact that $\frac{\eta_{t}}{H_{t}}\leqslant\frac{\alpha}{M}$ and the $M$-smoothness of $L$, we obtain
$\eta_{t}\left(L(\theta_{t+1})\!-\!L(\theta_{t})\!-\!\langle\nabla L(\theta_{t}),\theta_{t+1}\!-\!\theta_{t}\rangle\right)\!-\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\leqslant\frac{\eta_{t}M}{2}\|\theta_{t+1}\!-\!\theta_{t}\|_{2}^{2}-\!\frac{H_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}=0.$ (22)
By the convexity of $L$ and the Cauchy–Schwartz inequality,
$\eta_{t}(L(\theta_{t})-L(\theta_{t+1}))-\!\frac{\eta_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\leqslant-\eta_{t}\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle-\!\frac{\eta_{t}\alpha}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$
$\leqslant\frac{\eta_{t}M^{2}}{2\alpha}.$ (23)
Combining (21), (F) and (F), we conclude that
$H_{t}G_{t}\leqslant H_{t-1}G_{t-1}+\eta_{t}\langle\tilde{g}_{t}-\nabla L(\theta_{t}),\theta^{*}-\theta_{t+1}\rangle+\frac{\eta_{t}M^{2}}{2\alpha}$
as desired. ∎
Appendix G Proof of Corollary 1
Proof:
When $\frac{\eta_{t}}{H_{t}}=\frac{\alpha}{M}$, we write $G_{t}=\frac{H_{1}}{H_{t}}G_{1}+\frac{\sum_{k=2}^{t}E_{k}}{H_{t}}$ as
$G_{t}$ $=\frac{H_{1}}{H_{2}}\frac{H_{2}}{H_{3}}\dotsm\frac{H_{t-1}}{H_{t}}G_{1}+\frac{\sum_{k=2}^{t}E_{k}}{H_{t}}$
$=\Big{(}1-\frac{\eta_{2}}{H_{2}}\Big{)}\Big{(}1-\frac{\eta_{3}}{H_{3}}\Big{)}\dotsm\Big{(}1-\frac{\eta_{t}}{H_{t}}\Big{)}G_{1}+\frac{\sum_{k=2}^{t}E_{k}}{H_{t}}$
$=\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}G_{1}+\frac{\sum_{k=2}^{t}E_{k}}{H_{t}}$
From Lemmas 3 and 5, it follows that
$G_{t}=\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}\sum_{k=1}^{t}\eta_{k}}{2\alpha H_{t}}\\ +\frac{\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle}{H_{t}}$
By (20), we obtain the following sequence of upper bounds on the expectation of the final term above:
$\mathbb{E}\bigg{[}\frac{\sum_{k=1}^{t}\eta_{k}\langle\tilde{g}_{k}-\nabla L(\theta_{k}),\theta^{*}-\theta_{k+1}\rangle}{H_{t}}\bigg{]}$
$=\frac{R_{\Theta}C(t)}{m}\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}+(M+1)\sqrt{p}\sigma R_{\Theta}$
$=\frac{R_{\Theta}C(t)}{m}\sqrt{\frac{{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}^{2}}}+(M+1)\sqrt{p}\sigma R_{\Theta}$
$=\frac{R_{\Theta}C(t)\alpha}{mM}\sqrt{\frac{\sum_{k=1}^{t}H_{k}^{2}}{H_{t}^{2}}}+(M+1)\sqrt{p}\sigma R_{\Theta}$
$=\frac{R_{\Theta}C(t)\alpha}{mM}\sqrt{\Big{(}\frac{H_{1}}{H_{t}}\Big{)}^{2}+\ldots+\Big{(}\frac{H_{t}}{H_{t}}\Big{)}^{2}}+(M+1)\sqrt{p}\sigma R_{\Theta}$
$\leqslant\frac{R_{\Theta}C(t)\alpha}{mM}\sqrt{\frac{1-(1-\frac{\alpha}{M})^{2(t-1)}}{1-(1-\frac{\alpha}{M})^{2}}}+(M+1)\sqrt{p}\sigma R_{\Theta}$
$\leqslant\frac{R_{\Theta}C(t)}{m}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
Therefore, we conclude that:
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}\\ +\frac{R_{\Theta}C(t)}{m}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
When $\eta_{k}=\frac{2}{k+1}H_{k}$, the result of Case 2 follows by substituting this new choice of stepsize into Theorem 2. ∎
Appendix H Proof of Lemma 6
Proof:
We obtain the following inequality by substituting $C(t)=mt^{r}$ into the general bound for the smooth and strongly convex case in (12)
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\Big{(}1-\frac{\alpha}{M}\Big{)}^{t-1}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}\\ +R_{\Theta}t^{r}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
From the fact that $\|\theta^{*}-\theta_{0}\|_{2}^{2}\leqslant R_{\Theta}^{2}$, we can simplify the above upper bound of $\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]$ as follows
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\frac{\alpha R_{\Theta}^{2}}{2}e^{-\frac{\alpha}{M}(t-1)}+\frac{M^{2}}{2\alpha}\\ +R_{\Theta}t^{r}+(M+1)\sqrt{p}\sigma R_{\Theta}.$ (24)
We take the derivative of the upper bound in (24) with respect to $t$ and set the derivative to be zero as follows:
$-\frac{\alpha^{2}R_{\Theta}^{2}}{2M}e^{-\frac{\alpha}{M}(t-1)}+\frac{r\cdot R_{\Theta}}{t^{1-r}}=0.$ (25)
Then, we can rewrite (25) as
$te^{-\frac{\alpha}{(1-r)M}t}$ $=\left(\frac{2Mr}{\alpha^{2}R_{\Theta}\exp(\alpha/M)}\right)^{\frac{1}{1-r}}.$
Let $x=-\frac{\alpha}{(1-r)M}t$, we have
$xe^{x}=-\frac{\alpha}{(1-r)M}\left(\frac{2Mr}{\alpha^{2}R_{\Theta}\exp(\alpha/M)}\right)^{\frac{1}{1-r}}=:B(r).$
Therefore, we can obtain the solution $t_{0}$ by introducing the Lambert $W$ function $W_{-1}$ as follows:
$t_{0}=\left\lceil-\frac{(1-r)M}{\alpha}W_{-1}\left(B(r)\right)\right\rceil.$
Appendix I Proof of Theorem 3
Proof:
Substituting $C(t)=mt^{r}$ into (12), we obtain
$\mathbb{E}[L(\hat{\theta}_{t})-L(\theta^{*})]\leqslant\Big{(}\prod_{k=2}^{t}(1-\frac{\eta_{k}}{H_{k}})\Big{)}\frac{\alpha}{2}\|\theta^{*}-\theta_{0}\|_{2}^{2}+\frac{M^{2}}{2\alpha}\\ +R_{\Theta}t^{r}\cdot\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}+(M+1)\sqrt{p}\sigma R_{\Theta}.$
We now decompose the factor $\prod_{k=2}^{t}(1-\frac{\eta_{k}}{H_{k}})$ as follows:
$\prod_{k=2}^{t}(1-\frac{\eta_{k}}{H_{k}})$ $=\bigg{(}\prod_{k=2}^{t_{0}}(1-\frac{\alpha}{M})\bigg{)}\cdot\prod_{k=t_{0}+1}^{t}\Big{(}1-\frac{2}{k+1}\Big{)}$
$=\Big{(}1-\frac{\alpha}{M}\Big{)}^{t_{0}-1}\cdot\frac{t_{0}(t_{0}+1)}{t(t+1)}.$
Next, we decompose the term $\frac{1}{H_{t}}\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}$ as follows:
$\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}\!=\!\sqrt{\sum_{i=1}^{t_{0}}\Big{(}\frac{\eta_{i}}{H_{t}}\Big{)}^{2}+\sum_{i=t_{0}+1}^{t}\Big{(}\frac{\eta_{i}}{H_{t}}\Big{)}^{2}}.$ (26)
We now seek to bound $\sum_{i=t_{0}+1}^{t}(\frac{\eta_{i}}{H_{t}})^{2}$ in (26). We first derive the ratio $(\frac{\eta_{i-1}}{H_{t}})^{2}/(\frac{\eta_{i}}{H_{t}})^{2}$ for all $i\in\{t_{0}+2,\ldots,t\}$. Noting that $\frac{\eta_{i}}{H_{i}}=\frac{2}{i+1}$,
$\Big{(}\frac{\eta_{i-1}}{H_{t}}\Big{)}^{2}/\Big{(}\frac{\eta_{i}}{H_{t}}\Big{)}^{2}=\Big{(}\frac{\eta_{i-1}}{\eta_{i}}\Big{)}^{2}$
$=\Big{(}\frac{i+1}{i}\cdot\frac{H_{i-1}}{H_{i}}\Big{)}^{2}=\Big{(}\frac{i+1}{i}\cdot\frac{H_{i}-\eta_{i}}{H_{i}}\Big{)}^{2}$
$=\Big{(}\frac{i+1}{i}\cdot\left(1-\frac{\eta_{i}}{H_{i}}\right)\Big{)}^{2}=\Big{(}\frac{i+1}{i}\cdot\Big{(}1-\frac{2}{i+1}\Big{)}\Big{)}^{2}$
$=\Big{(}\frac{i-1}{i}\Big{)}^{2}.$
Hence, we have
$\sum_{i=t_{0}+1}^{t}\Big{(}\frac{\eta_{i}}{H_{t}}\Big{)}^{2}$
$=\Big{(}\frac{\eta_{t}}{H_{t}}\Big{)}^{2}\bigg{(}1\!+\!\Big{(}\frac{t-1}{t}\Big{)}^{2}\!+\!\Big{(}\frac{t-2}{t}\Big{)}^{2}\!+\!\dotsm\!+\!\Big{(}\frac{t_{0}+1}{t}\Big{)}^{2}\bigg{)}$
$=\Big{(}\frac{\eta_{t}}{H_{t}}\Big{)}^{2}\bigg{(}\frac{t^{2}+(t-1)^{2}+(t-2)^{2}+\dotsm+(t_{0}+1)^{2}}{t^{2}}\bigg{)}$
$=\Big{(}\frac{2}{t+1}\Big{)}^{2}\bigg{(}\frac{\sum_{i=t_{0}+1}^{t}i^{2}}{t^{2}}\bigg{)}.$ (27)
Next, we can bound the remaining part in (26) as follows
$\sum_{i=1}^{t_{0}}\Big{(}\frac{\eta_{i}}{H_{t}}\Big{)}^{2}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\frac{(\frac{\eta_{t_{0}}}{H_{t}})^{2}(1-(1-\alpha/M)^{2t_{0}})}{1-(1-\alpha/M)^{2}}$
$\leqslant\frac{(\frac{\eta_{t_{0}}}{H_{t}})^{2}}{1-(1-\alpha/M)^{2}}\stackrel{{\scriptstyle(b)}}{{\leqslant}}\frac{M^{2}}{\alpha^{2}}\Big{(}\frac{\eta_{t_{0}}}{H_{t}}\Big{)}^{2},$
where $(a)$ is due to the fact that $(\frac{\eta_{j}}{H_{t}})^{2}/(\frac{\eta_{j+1}}{H_{t}})^{2}=(1-\alpha/M)^{2}$ for all $j\in\{1,\ldots,t_{0}-1\}$ and $(b)$ is due to the fact that $1-(1-\alpha/M)^{2}\geqslant\alpha^{2}/M^{2}$ when $\alpha/M\leqslant 1$.
To compute $(\frac{\eta_{t_{0}}}{H_{t}})^{2}$, consider
$\frac{\eta_{t_{0}+1}}{H_{t_{0}+1}}$ $=\frac{2}{t_{0}+2},$ $\quad\frac{\eta_{t_{0}+1}}{H_{t_{0}}+\eta_{t_{0}+1}}$ $=\frac{2}{t_{0}+2},$
$\frac{H_{t_{0}}}{\eta_{t_{0}+1}}$ $=\frac{t_{0}}{2},$ $\quad\frac{\eta_{t_{0}}}{\eta_{t_{0}+1}}$ $=\frac{\alpha}{M}\cdot\frac{t_{0}}{2},$
where the final equality is because $\eta_{t_{0}}/H_{t_{0}}=\alpha/M$.
From $\big{(}\frac{\eta_{t_{0}+1}}{H_{t}}\big{)}^{2}=\big{(}\frac{\eta_{t}}{H_{t}}\big{)}^{2}\big{(}\frac{t_{0}+1}{t}\big{)}^{2}$, we have the equality $\big{(}\frac{\eta_{t_{0}}}{H_{t}}\big{)}^{2}=\frac{\alpha^{2}}{M^{2}}\big{(}\frac{t_{0}(t_{0}+1)}{t(t+1)}\big{)}^{2}$. Hence,
$\sum_{i=1}^{t_{0}}\Big{(}\frac{\eta_{i}}{H_{t}}\Big{)}^{2}\leqslant\Big{(}\frac{t_{0}(t_{0}+1)}{t(t+1)}\Big{)}^{2}.$ (28)
By substituting (27) and (28) back into (26), we obtain
$\frac{\sqrt{\sum_{k=1}^{t}\eta_{k}^{2}}}{H_{t}}\!$ $\leqslant\!\sqrt{\Big{(}\frac{t_{0}(t_{0}+1)}{t(t+1)}\Big{)}^{2}\!+\!\Big{(}\frac{2}{t(t+1)}\Big{)}^{2}\sum_{k=t_{0}+1}^{t}k^{2}}$
completing the proof upon using the definition of $A(t,t_{0})$. ∎